Describe how the x- and y-coordinates of the given quadratic functions differ from the x- and y-coordinates of corresponding points of y=x^2

y=-2(x-4)^2
1
Expert's answer
2020-12-11T02:52:18-0500

y =-2(x-4)2

Direction: directed downwards

Top: (4.0)

Focus: (4, −1 / 8).

Symmetry axis: x = 4

Guide: y = 1/8



y = x2

Direction: directed upwards

Top: (0.0)

Focus: (0.1 / 4).

Symmetry axis: x = 0

Guide: y = -1/4